Using identities, evaluate
10.5 x 9.5 ?????????????????

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

10.5 x 9.5
= (10 + 0.5)(10 - 0.5)
= 10^2- 0.5^2
= 100 - 0.25
= 99.75

Thanks